OpenStudy (anonymous):

please help thankxx Which value of θ satisfies the equation 2cos^2 θ - cosθ=0

OpenStudy (mertsj):

\[\cos \theta(2\cos \theta-1)=0\]

OpenStudy (anonymous):

thankxs

OpenStudy (anonymous):

wait hold on a min.

OpenStudy (anonymous):

i have a question

OpenStudy (mertsj):

\[\cos \theta=0 or \cos \theta=\frac{1}{2}\]

OpenStudy (mertsj):

\[\theta=\frac{pi}{2}, \frac{3pi}{2} + 2pin\]

OpenStudy (anonymous):

this a multipe chioice question and ur answer isnt there

OpenStudy (mertsj):

It probably is and you don't recognize it. Anyway I have not completed the question. There are more answers. Why don't you make my life easier by posting the answers?

OpenStudy (anonymous):

nm i got it

OpenStudy (mertsj):

\[\theta=\frac{pi}{3}, \frac{5pi}{3}+2pin\]

OpenStudy (anonymous):

i got it thank u